OpenStudy (anonymous):

7(2+3)-3+____-8(6-9+2)-3=32

OpenStudy (anonymous):

7(5)-3+x-8(-1)-3=32 35-3+x+8-3=32 32+x+5=32 37+x=32 x=-5